OpenStudy (anonymous):

A certain tank can be filled by two pipes in 80 minutes. The larger pipe by itself can fill the tank in 2 hours less than the smaller pipe by itself. How long does each pipe take to fill the tank on its own?

OpenStudy (anonymous):

80 minutes = 1 1/3 hour

OpenStudy (lgbasallote):

actually it's a work problem where 1/x + 1/y = 1/t where t = total time working together

OpenStudy (anonymous):

ooh yeah. . my bad. .

OpenStudy (lgbasallote):

i tried solving it...my brain hurt =)))

OpenStudy (anonymous):

let x= smaller pipe y= bigger pipe so, equations will be, (1) 1/x +1/y = 1 1/3 (2) y = x-2

OpenStudy (lgbasallote):

actualy 1/x + 1/y + 1/4/3

OpenStudy (anonymous):

i think it's 4/3 or 1 1/3, not 1/4/3

OpenStudy (dumbcow):

\[\frac{1}{x}+\frac{1}{x-2} = \frac{3}{4}\] \[\frac{2(x-1)}{x(x-2)} = \frac{3}{4}\] \[8x-8 = 3x^{2}-6x\] \[3x^{2} -14x+8 = 0\] \[(3x-2)(x-4) = 0\] x = 4 so it takes larger pipe 2 hrs and smaller pipe 4 hrs

OpenStudy (anonymous):

That's what the book says, how did you get from step 1 to step 2

OpenStudy (anonymous):

Don't worry I got it now

OpenStudy (dumbcow):

combined fractions....ok:)

OpenStudy (anonymous):

In step one, why is it equal to 3/4 shouldn't it be 4/3

OpenStudy (dumbcow):

each fraction represents the rate it takes to fill up the tank, T/hr it took 4/3 hr so rate is 1/(4/3) = 3/4 tanks per hr

OpenStudy (anonymous):

Ahh... Thanks

OpenStudy (dumbcow):

welcome